OpenStudy (anonymous):

Mr. chow drives 112.5 miles at 45 miles per hour 357.5 miles at 55 miles per hour and 195 miles at 65 miles per hour how many hours did he spend driving altogether

OpenStudy (anonymous):

time is distance divided by rate divide each, then add them up

OpenStudy (anonymous):

\[\huge t=\frac{ d }{ r }\]

OpenStudy (anonymous):

\[\huge hours=\frac{ 112.5 }{ 45 }+\frac{ 357.5 }{ 55 }+\frac{ 195 }{ 65}\]

OpenStudy (anonymous):

\[\huge hours= 2.5 + 6.5 + 3\]Add them together.